OpenStudy (anonymous):

find the indicated limit:

OpenStudy (anonymous):

lim x->2+ \[\frac{ 7\sqrt{x-2^{3}} }{ x-2 }\]

myininaya (myininaya):

Your function isn't even defined at or around 2=x since x-8>=0 when x>=8 and 2 is definitely no in that interval

OpenStudy (anonymous):

i dont get it

myininaya (myininaya):

You have x-8 inside that square root

myininaya (myininaya):

sqrt(x-8) exists over the reals when x-8>=0

geerky42 (geerky42):

Domain for that expression is \([8,\infty)\) There is nothing at around x=2

myininaya (myininaya):

You could have meant a different function though. Look at what you wrote and confirm that is or isn't the function we should be looking at.